# Electromagnetic Flow Control in Hypersonic Rarefied Environment

### v1

@@ 2 UGKWP Method
The UGKWP method is used to perform numerical simulations in this study. In the framework of finite volume method, the cell-averaged macroscopic conservative variable $\boldsymbol{W}_{\alpha}=(\rho_{\alpha},\rho_{\alpha}\boldsymbol{U}_{\alpha},\rho_{\alpha}\mathscr{E}_{\alpha})$, i.e, mass, momentum and energy densities of species $\alpha$ on a physical cell $\Omega_{i}$ are updated by the flux through cell interfaces, electromagnetic forces and cross-species collisions:
$(\boldsymbol{W}_{\alpha})_{i}^{n+1}=(\boldsymbol{W}_{\alpha})_{i}^{n}-\frac{\Delta t}{|\Omega_{i}|}\sum_{s\in\partial\Omega_{i}}|l_{s}|(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}+\frac{\Delta t}{\tau_{\alpha}}(\bar{(\boldsymbol{W}}_{\alpha})_{i}^{n}-(\boldsymbol{W}_{\alpha})_{i}^{n})+\Delta t(\boldsymbol{S}_{\alpha})_{i}^{n+1},$ (1)
@@
$(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\boldsymbol{u}\cdot\boldsymbol{n}_{s}f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)\boldsymbol{\Psi}\mathrm{d}\boldsymbol{\Xi}\mathrm{d}t,$ (2)
@@
where $\boldsymbol{\Psi}=(1,\boldsymbol{u},\frac{1}{2}(\boldsymbol{u}^{2}+\boldsymbol{\xi}^{2}))$ is the conservative moments of distribution functions with $\boldsymbol{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{n})$ the internal degree of freedom. $\mathrm{d}\boldsymbol{\Xi}=\mathrm{d}\boldsymbol{u}d\boldsymbol{\xi}$ is the volume element in the phase space. For cell interface $l_{s}$, the distribution function $f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)$ cannot be analytically represented in the transitional and collisionless regimes, thus requiring direct tracking of its evolution. The evolution of the distribution function in non-continuum regimes is crucial for realizing multiscale transport in the UGKWP method, as detailed in [23]. $(\bar{\boldsymbol{W}}_{\alpha})_{i}=((\rho_{\alpha})_{i},(\rho_{\alpha}\bar{\boldsymbol{U}}_{\alpha})_{i},(\rho_{\alpha}\bar{\mathscr{E}}_{\alpha})_{i})$ where $(\bar{\boldsymbol{U}}_{\alpha})_{i}$ and (¯$(\bar{\mathscr{E}}_{\alpha})_{i}$ are post-collision velocity and energy in multispecies kinetic model. In this paper, the Andries-Aoki-Perthame (AAP) model is used to model cross-species collisions. In the AAP model, post-collision temperature and velocity are calculated as:
@@
$\bar{\boldsymbol{U}}_{\alpha}$ $=\boldsymbol{U}_{\alpha}+\frac{\tau_{\alpha}}{m_{\alpha}}\sum_{k=1}^{N}2\mu_{\alpha k}\chi_{\alpha k}n_{k}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right),$ (3)
@@
$\bar{T}_{\alpha}$ $=T_{\alpha}-\frac{m_{\alpha}}{3k_{B}}\left(\boldsymbol{U}_{\alpha}^{*}-\boldsymbol{U}_{\alpha}\right)^{2}+\tau_{\alpha}\sum_{k=1}^{N}\frac{4\mu_{\alpha k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\left(T_{k}-T_{\alpha}+\frac{m_{k}}{3k_{B}}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right)^{2}\right),$
where $\mu_{\alpha k}=m_{\alpha}m_{k}/(m_{\alpha}+m_{k})$ is reduced mass, mean relaxation time $\tau_{\alpha}$ is determined by $1/\tau_{\alpha}=\sum_{k=1}^{m}\chi_{\alpha k}n_{k}$, and interaction coefficient $\chi_{\alpha k}$ for hard sphere model is [24]:
$\chi_{\alpha k}=\frac{4\sqrt{\pi}}{3}\left(\frac{2k_{B}T_{\alpha}}{m_{\alpha}}+\frac{2k_{B}T_{k}}{m_{k}}\right)^{1/2}\left(\frac{d_{\alpha}+d_{k}}{2}\right)^{2},$
where $d_{\alpha},d_{k}$ are the diameters of the particles and can be approximated by
$(d_{\alpha}+d_{k})^{2}=\frac{1}{\sqrt{2}\pi(n_{\alpha}+n_{k})\text{Kn}\text{L}},$
where Kn is Knudsen number, L is reference length. The molecular diameter is modified by the temperature as
$\left(\frac{d}{d_{ref}}\right)=\left(\frac{T}{T_{ref}}\right)^{-\xi/2},$
where $\xi=\omega-\frac{1}{2}$, and $\omega$ is the power coefficient between viscosity and temperature[25]. In this work, $\omega=0.81$ for argon atom. The cross section between electron and neutral particles is described by a fitted curve from Devoto [26].
In this study, the low magnetic Reynolds number approximation is employed, which implies that the externally applied magnetic field dominates over the induced magnetic field generated by the current. This assumption holds when the electrical conductivity is sufficiently low. Additionally, an approximate Debye length is implemented to alleviate numerical stiffness, as resolving the exact Debye length would be computationally prohibitive, particularly in three-dimensional simulations.
@@ 3 Numerical Studies
3.1 Hypersonic Flow Around a Sphere
[…]
@@ 3.1 Hypersonic Flow Around a Sphere
Figure 1 displays the contour distributions of density, velocity, temperature, and local Knudsen number on and around the wall. The significant variation in Knudsen number, from 0.031 to values greater than 1 across the entire flow field (Figure 1(d)), highlights the algorithm’s effectiveness in capturing flow physics across different rarefied flow regimes. The surface quantities are presented in Figure 2. The present results exhibit excellent agreement with the reference data generated by the Unified Gas-Kinetic Scheme (UGKS) code[27]. This shows the accuracy of the current solver in the ionization-free condition.
Figure 1: Hypersonic flow around a sphere at Ma=4.25 and Kn=0.031 by the UGKWP method. a) density, b) velocity in x direction, c) temperature, and d) local Knudsen number contours.
@@ 3.2 Hypersonic Pre-ionized Argon Flow Around a Hemisphere
To further assess the code’s accuracy, we conduct a three-dimensional simulation of Mach 4.75 argon flow around a hemisphere, replicating the experimental conditions reported by Kranc et al. [6]. The hemisphere radius is defined as $r_{n}=0.01905$ m. Figure 3(a) illustrates the computational domain geometry employed in this study. The simulation imposes a hypersonic freestream flow in the $-x$ direction incident upon the hemisphere. We generate an unstructured hexahedral mesh to discretize the computational domain surrounding the hemisphere.
(a) Computational domain geometry
(b) Dipole magnetic field contours ($B_{\text{max}}=0.387\ \text{T}$).
@@
where, $\hat{x}=x/r_{n}$, $\hat{y}=y/r_{n}$, and $\hat{z}=z/r_{n}$ represent the coordinates normalized by the hemisphere radius, $r_{n}$. The parameter $B_{max}$ denotes the maximum magnetic field strength, which is attained at the stagnation point. The magnetic field magnitude distribution is depicted in Figure 3(b). In the following analysis, we employ $B_{max}$ to characterize the magnetic field strength.
Simulations are performed at four distinct magnetic field strengths: 0.223 T, 0.316 T, 0.387 T, and 0.447 T. The simulations employ a two-step approach. Initially, the flow field is computed in the absence of a magnetic field until a steady-state solution is achieved. Subsequently, the magnetic field is introduced, and the simulation is continued until a new steady-state solution, reflecting the influence of the applied magnetic field, is obtained.
Figure 4: Contour plots of (a) density, (b) x-velocity, (x) temperature, and (d) heat flux, comparing conditions with and without an applied magnetic field. The upper half of each plot shows results for $B_{0}=0T$, while the lower half illustrates the effects of a $B_{0}=0.387T$ background magnetic field. The data clearly demonstrate that the magnetic field increases the shock standoff distance and markedly decreases the heat flux.
Figure 5: Contour plots on the wall of heat flux of (a) Ar, (b) Ar+, (c) electron, comparing conditions with and without an applied magnetic field. The upper half of each plot shows results for $B_{0}=0T$, while the lower half illustrates the effects of a $B_{0}=0.387T$ background magnetic field. The data clearly demonstrate that the magnetic field decreases the heat flux of Ar, but increases that of Ar+ and electrons.
Kranc et al. [6] observed an increase in shock standoff distance when an external magnetic field was applied. This increase was determined by comparing photographic images of the flow with and without the magnetic field, with the upstream edge of the shock inferred from the boundary of the flow’s luminosity. In computational simulations, a universally accepted definition of the shock edge remains elusive. Bisek et al. [28] defined the shock location as the point along the stagnation line where the density ratio exceeds the theoretical infinite Mach number threshold for a normal shock wave in an ideal gas. However, this density ratio proves excessively high for rarefied flow conditions. In the present study, we estimate the shock edge location to occur where the density ratio exceeds twice the freestream density, a criterion that aligns with the shock edge observed in density contours.
Contour distributions of density, velocity, temperature, and heat flux on and around the wall are presented in Figures 4 and 5. The upper half of each plot represents the results without a magnetic field ($B_{0}=0\text{ T}$), and the lower half shows the results of a $B_{0}=0.387\text{ T}$ background magnetic field. The magnetic field clearly increases the shock standoff distance. For argon, a smaller temperature gradient leads to a reduction in heat flux. Conversely, the heat flux for argon ions ($\text{Ar}^{+}$) and electrons ($\text{e}^{-}$) increases, possibly due to direct magnetic heating. However, the overall heat flux is lower due to the low concentration of $\text{Ar}^{+}$ and $\text{e}^{-}$ relative to Ar, and the minor increment in heat flux is trivial compared to the decrement observed in Ar.
Figure 6 depicts the variation in shock standoff distance with changing argon ion-electron mass ratios in our simulations. We observe a gradual convergence of the standoff distance as the mass ratio increases. Critically, our calculations, performed with the multiscale solver UGKWP to accurately capture the Ar flow field, show better agreement with the experimental results of Kranc et al. This contrasts with reference computational results from MHD employing different conductivity models. The findings clearly highlights that an accurate Ar flow field is fundamental to accurately predicting the shock standoff distance.
@@
Figure 6: Changes in shock standoff distance as a function of magnetic field strength for Mach 4.75 pre-ionized argon flow around a hemispherical geometry, parameterized by mass ratio. Experimental data are from Kranc et al. [6], while computational results using MHD methods are from Bisek et al. [28] using various conductivity model. (Experimental uncertainties is ±$\pm 10\%$)

### v2

@@ 3. Step 3:
Based on the finite volume framework, the cell-averaged macroscopic conservative variables on a physical cell $\Omega_{i}$ are updated by the numerical flux through cell interfaces, electromagnetic forces and cross-species collisions:
$(\boldsymbol{W}_{\alpha})_{i}^{n+1}=(\boldsymbol{W}_{\alpha})_{i}^{n}-\frac{\Delta t}{|\Omega_{i}|}\sum_{s\in\partial\Omega_{i}}|l_{s}|(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s},$
@@
$(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\boldsymbol{u}\cdot\boldsymbol{n}_{s}f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)\boldsymbol{\Psi}\mathrm{d}\boldsymbol{\Xi}\mathrm{d}t,$
@@
where $\boldsymbol{\Psi}=(1,\boldsymbol{u},\frac{1}{2}(\boldsymbol{u}^{2}+\boldsymbol{\xi}^{2}))$ is the conservative moments of distribution functions with $\boldsymbol{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{n})$ the internal degree of freedom. $\mathrm{d}\boldsymbol{\Xi}=\mathrm{d}\boldsymbol{u}d\boldsymbol{\xi}$ is the volume element in the phase space. For the cell interface $l_{s}$, the distribution function $f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)$ cannot be analytically represented in the transitional and collisionless regimes, thus requiring direct tracking of its evolution as introduced in the previous paragraph. This numerical flux of the macroscopic conservative variable can be split into the equilibrium flux and free streaming flux according to Eq.(7). The equilibrium flux is
@@
$(\mathscr{F}_{\boldsymbol{W}})_{s}^{g}=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\boldsymbol{u}\cdot\boldsymbol{n}_{s}\left[\frac{1}{\tau}\int_{0}^{t}g(\boldsymbol{x}^{\prime},\boldsymbol{u},\boldsymbol{\xi},t^{{}^{\prime}})e^{-\left(t-t^{\prime}\right)/\tau}\mathrm{d}t^{\prime}\right]\boldsymbol{\Psi}\mathrm{d}\boldsymbol{\Xi}\mathrm{d}t,$ (10)
and the free streaming flux is
@@
$(\mathscr{F}_{\boldsymbol{W}})_{s}^{f}=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\boldsymbol{u}\cdot\boldsymbol{n}_{s}\left[e^{-t/\tau}f_{0}\left(\boldsymbol{x}-\boldsymbol{u}t\right)\right]\boldsymbol{\Psi}\mathrm{d}\boldsymbol{\Xi}\mathrm{d}t.$ (11)
The equilibrium flux can be calculated directly from the macroscopic conservative variables. Assume in the equation (2.2), $\boldsymbol{x}_{s}=\boldsymbol{0}$ and $t^{n}=0$, the equilibrium $g$ at the cell interface is obtained by conservation constraint
$\int{g\boldsymbol{\Psi}}d\boldsymbol{\Xi}=\int_{\boldsymbol{v}\cdot\boldsymbol{n}>0}{g^{l}\boldsymbol{\Psi}}d\boldsymbol{\Xi}+\int_{\boldsymbol{v}\cdot\boldsymbol{n}<0}{g^{r}\boldsymbol{\Psi}}d\boldsymbol{\Xi}.$
The spatial and time derivatives can be obtained,
$\int{g_{x}\boldsymbol{\Psi}}d\boldsymbol{\Xi}=\boldsymbol{W}_{x},\quad\int{g_{t}\boldsymbol{\Psi}}d\boldsymbol{\Xi}=-\int\boldsymbol{u}\cdot g_{x}\boldsymbol{\Psi}d\boldsymbol{\Xi},$
where $g^{l}$ and $g^{r}$ are the equilibrium distributions according to the reconstructed left and right side conservative variables at cell interface $\boldsymbol{W}^{l},{\ \boldsymbol{W}}^{r}$, and $\boldsymbol{W}_{x}$ is the reconstructed spatial derivative of conservative variables at cell interface. The van Leer limiter is used to achieve a second-order accuracy in space reconstruction. Substituting the reconstructed equilibrium distribution into the equilibrium flux, we have
$(\mathscr{F}_{\boldsymbol{W}})_{s}^{g}=\int\boldsymbol{u}\cdot\boldsymbol{n}_{s}\left(C_{1}g_{0}+C_{2}\boldsymbol{u}\cdot{\boldsymbol{g}}_{0x}+C_{3}g_{0t}\right)\Psi d\boldsymbol{\Xi},$
where the time integration coefficients are
$C_{1}=\Delta t-\tau\left(1-e^{-\Delta t\text{/}\tau}\right),$
$C_{2}=2\tau^{2}(1-e^{-\Delta t\text{/}\tau})-\tau\Delta t-\tau\Delta te^{-\Delta t\text{/}\tau},$
$C_{3}=\frac{\Delta t^{2}}{2}-\tau\Delta t+\tau^{2}(1-e^{\Delta t/\tau}).$
Next we consider the free stream flux $(\mathscr{F}_{\boldsymbol{W}})_{s}^{f}$ . As stated in the last subsection, the initial distribution is represented partially by an analytical distribution $g_{a}^{+,c}$, and partially by particles, and therefore the free stream flux is also calculated partially from the reconstructed analytical distribution as $(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,w}$, and partially from particles as $(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,p}$. The initial analytical distribution $g_{\alpha}^{+,c}$ is reconstructed as
$g_{0}^{+,c}\left(\boldsymbol{x},\boldsymbol{u}\right)=g_{0}^{+,c}+{\boldsymbol{g}}_{0x}^{+,c}\cdot\boldsymbol{x},$
which gives
$(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,w}=\int\boldsymbol{u}\cdot\boldsymbol{n}_{s}\left(C_{4}g_{0}^{+,c}+C_{5}\boldsymbol{u}\cdot{g}_{0x}^{+,c}\right)\boldsymbol{\Psi}d\boldsymbol{\Xi},$
where the time integration coefficients are
$C_{4}=\tau\left(1-e^{-\Delta t\text{/}\tau}\right)-\Delta te^{-\Delta t\text{/}\tau},,$
$C_{5}=\tau\Delta te^{-\Delta t\text{/}\tau}-\tau^{2}\left(1-e^{-\Delta t\text{/}\tau}\right)+\frac{\Delta t^{2}}{2}e^{-\Delta t\text{/}\tau}.$
The net particle flux $(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,p}$ is calculated as
$(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,p}=\sum_{k\in P_{\partial\Omega_{i}^{+}}}\boldsymbol{W}_{P_{k}}-\sum_{k\in P_{\partial\Omega_{i}^{-}}}\boldsymbol{W}_{P_{k}},$
where $\boldsymbol{W}_{P_{k}}=\left(m_{k},m_{k}\boldsymbol{v}_{k},\frac{1}{2}m_{k}\boldsymbol{v}_{\boldsymbol{k,}}^{2}\right),P_{\partial\Omega_{i}^{-}}$ is the index set of the particles streaming out of cell $\Omega_{i}$ during a time step, and $P_{\partial\Omega_{i}^{+}}$ is the index set of the particles streaming into cell $\Omega_{i}$. Finally, the update formula for conservative variables is
$\begin{split}\boldsymbol{W}_{\boldsymbol{i}}^{*}=\boldsymbol{W}_{\boldsymbol{i}}^{\boldsymbol{n}}-&\sum_{s}{\frac{\Delta t}{\left|\Omega_{i}\right|}\left|l_{s}\right|(\mathscr{F}_{\boldsymbol{W}})_{s}^{g}}-\sum_{s}{\frac{\Delta t}{\left|\Omega_{i}\right|}\left|l_{s}\right|(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,w}}+\frac{1}{\left|\Omega_{i}\right|}(\mathscr{F}_{\boldsymbol{W}})_{s}^{f,p}\end{split}$ (12)
The evolution of the macroscopic conservative variables due to species transport and diffusion has been solved by the coupled evolution of the microscopic particle transport and macroscopic fluid dynamics.
Subsequently, we solve the $\mathcal{L}_{2}$. The moment equation of Eq. 4, is first solved implicitly. Particles are then accelerated using the updated $\boldsymbol{E}^{*}$. Further details regarding the moment equation and the implicit linear system can be found in [24]. Next, we solve $\mathcal{L}_{3}$, the moment equation of Eq.(5) is a set of nonlinear ordinary equations:
$\frac{\partial\boldsymbol{U}_{\alpha}}{\partial t}=\frac{1}{m_{\alpha}}\sum_{k=1}^{N}2\mu_{\alpha k}\chi_{\alpha k}n_{k}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right),$ (13)
$\frac{\partial E_{\alpha}}{\partial t}=\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\boldsymbol{U}_{\alpha}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right)+\frac{3k_{B}}{2m_{\alpha}}\sum_{k=1}^{N}\frac{4\mu_{\alpha k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\left(T_{k}-T_{\alpha}+\frac{m_{k}}{3k_{B}}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right)^{2}\right)$ (14)
Discretize with the backward Euler scheme and solve it with the Newton iteration solver. The details is in B. Finally, in $\mathcal{L}_{4}$, Eq.(6) is solved by the flux vector splitting method.
The overall algorithm flow chart is shown in Figure 1.
Figure 1: Flow chart of UGKWP method
@@ 3 Numerical Studies
3.1 Single Species Hypersonic Nonequilibrium Argon Flow Around a Sphere
To validate the computational code’s accuracy in simulating rarefied gas dynamics under conditions of no ionization, we perform a numerical simulation of a hypersonic neutral argon gas flow around a three-dimensional sphere. This is crucial as an accurate representation of the neutral gas flow field is essential for the subsequent modeling of weakly ionized plasma phenomena. The simulation parameters are set to a Mach number of 4.25 and a Knudsen number of 0.031, with the sphere diameter, $d=2$ m, serving as the characteristic length. The freestream quantities of argon atoms compute the characteristic Knudsen number. The computational domain is discretized using 132,480 hexahedral cells, with the first layer adjacent to the sphere surface having a height of 0.005dd. A comprehensive list of simulation parameters is provided in Table 1. Boundary conditions consist of a far-field inflow and outflow condition at the domain’s outer boundaries and an isothermal wall condition imposed on the sphere’s surface. The reference number of particles sampled per cell is 400.
@@ 3.1 Single Species Hypersonic Nonequilibrium Argon Flow Around a Sphere
Figure 2 displays the contour distributions of density, velocity, temperature, and local Knudsen number on and around the wall. The local Knudsen number is defined as $Kn_{local}=\frac{\lambda}{\rho/|\nabla\rho|}$, where $\lambda$ is local molecular mean free path. The significant variation in Knudsen number, from 0.031 to values greater than 1 across the entire flow field (Figure 2(d)), highlights the algorithm’s effectiveness in capturing flow physics across different rarefied flow regimes. The surface quantities are presented in Figure 3. The present results exhibit excellent agreement with the reference data generated by the Unified Gas-Kinetic Scheme (UGKS) code [29], which is well-validated by the Direct Simulation Monte Carlo (DSMC) method. This shows the accuracy of the current solver in the ionization-free condition.
Figure 2: Hypersonic flow around a sphere at Ma=4.25 and Kn=0.031 by the UGKWP method. a) density, b) Mach number, (c) temperature and (d) local Knudsen number.
@@ 3.2 Hypersonic Pre-ionized Argon Flow Around a Hemisphere
In this section, we conduct a three-dimensional simulation of Mach 4.75 argon flow around a hemisphere, replicating the experimental conditions reported by Kranc et al. [6]. To further verify the code’s accuracy, we first compare the results with the experiment’s results. Then we change the Knudsen number of argon atoms to see how the control effects vary in different rarefied degrees.
3.2.1 Validation against experiments
The hemisphere radius is defined as $r_{n}=0.01905$ m. The simulation imposes a hypersonic freestream flow in the x-x direction incident upon the hemisphere as shown in Figure 4(a). $1/4$ of the hemisphere is simulated as shown in Figure 4(b). The height of the first layer is 0.001m, and the total cell number is 5568.
(a) Case setting
(b) Mesh setting
(c) Magnetic field setting $\frac{|\boldsymbol{B}|}{|\boldsymbol{B}_{max}|}$.
@@ 3.2.1 Validation against experiments
where, $\hat{x}=x/r_{n}$, $\hat{y}=y/r_{n}$, and $\hat{z}=z/r_{n}$ represent the coordinates normalized by the hemisphere radius, $r_{n}$. The parameter $B_{max}$ denotes the maximum magnetic field strength, which is attained at the stagnation point. The distribution of magnetic field magnitude is depicted in Figure 4(c). In the following analysis, we employ $B_{max}$ to characterize the magnetic field strength. Simulations are performed at four distinct magnetic field strengths: 0.223 T, 0.316 T, 0.387 T, and 0.447 T.
In this case, ArAr-ArAr, $^{+}-Ar+Ar^{+}$, $^{-}-ee^{-}$ collision is modeled as VHS collision with $\omega=0.81$ and $d=4.17\times{10^{-10}}$m. $Ar^{+}, ArAr-$$e^{-}$, $Ar^{+}$-$e^{-}$ cross section is described by fitter curves from Devoto [31] as shown in Table 3.
Collision pair Fitted curve
$Ar+Ar^{+}$ $0.00182T_{Ar^{+}}+116.1$
ArAr-$e^{-}$ $0.0004T_{e^{-}}0.3838$
$Ar^{+}$-$e^{-}$ $0.001$
[…]
Electron mass was artificially increased ($m_{Ar^{+}}/m_{e^{-}}=10$) to be computationally tractable. This modification, however, reduces the electrical conductivity. Therefore, the magnetic field magnitude is artifially increased, following the scaling by the similarity parameter $Q=\frac{\sigma^{cond}|\boldsymbol{B}_{0}|^{2}L_{0}}{\rho_{0}|\boldsymbol{U}_{0}|}$, see more explanation in A.
Current formulation allows for the observation of distinct behaviors among atoms, ions, and electrons, including velocity and temperature slip arising from rarefied effects. Figures 5 and 6 display contour distributions, with upper halves showing simulations without an external magnetic field and lower halves depicting results with background magnetic field strengths ($B=0.447$ T). The data confirm that an applied magnetic field enhances shock standoff distance. Figures 5(b) and 5(c) demonstrate that after the application of a magnetic field, both ions and electrons are significantly trapped near the stagnation point. This trapping occurs due to the localized maximum strong magnetic field at this location. The collision coupling between these charged particles and argon atoms also contributes to a slight accumulation of argon atoms in the stagnation region, as depicted in Figure 5(a). Figures 5(d), 5(e), and 5(f) present the x-velocity of $Ar$, $Ar^{+}$, and $e^{-}$. In the stagnation region, $Ar^{+}$ velocity closely follows $Ar$ velocity. In the shoulder region  [^1] The shoulder region comprises arc degrees spanning from approximately $30^{\circ}$ to $60^{\circ}$, with the stagnation line serving as the $0^{\circ}$ reference., outside of shock, their velocities diverge due to the magnetic field influence on $Ar^{+}$. Electron velocity deviates significantly from heavy particles as the magnetic field more easily accelerates them, and the velocity decoupling is further promoted by low collision frequencies in rarefied environments. Figures 5(g) to 5(i) show similar decoupling in temperature.
The present formulation permits the observation of electric fields generated by diverse mechanisms, contrasting with MHD-based methods that primarily consider resistivity. Figures 6(a) depict the electric field magnitude with and without an external magnetic field. Without a magnetic field, the primary electric field is observed in the stagnation region, driven by a strong pressure gradient across the shock front that separates electrons and ions. The non-neutral region is amplified here by an increased Debye length. In reality, this non-neutrality is localized to the shock front and wall sheath. The application of a magnetic field induces an additional charge separation in the shoulder region due to the different responses of ions and electrons to the magnetic field. Figure 6(b) shows a strong current density concentrated in the stagnation region. The current density here is computed by $\boldsymbol{J}=q_{i}n_{i}\boldsymbol{U}_{i}+q_{e}n_{e}\boldsymbol{U}_{e}$. Figure 6(c) reveals a substantial Lorentz force near this area. The Lorentz force is computed as $\boldsymbol{F}=\boldsymbol{J}\times\boldsymbol{B}$. The force does not precisely align with the peak current density and magnetic field strength because the magnitude of the Lorentz force $\boldsymbol{J}\times\boldsymbol{B}$ is also critically dependent on the angle between the current density $\boldsymbol{J}$ and the magnetic field $\boldsymbol{B}$.
The dependence of shock standoff distance on applied magnetic field strength is presented in Figure 7. The shock front location is determined using the density jump ratio, which is defined as [30] :
$\lim_{M_{1}\rightarrow\infty}\frac{\rho_{2}}{\rho_{1}}=\frac{\gamma+1}{\gamma-1}$
Here, $M_{1}$ denotes the upstream Mach number, $\gamma$ represents the ratio of specific heats, $\rho_{1}$ is the upstream density, and $\rho_{2}$ is the downstream density. For argon gas, with $\gamma_{Ar}=5/3$, this density ratio is calculated to be 4. Our simulations exhibit good agreement with experimental data. In addition to experimental results, reference simulation data from Bisek et al. [30], employing three distinct conductivity models, are included for comparison. The primary distinction between the present formulation and Bisek’s work lies in our consideration of non-equilibrium transport phenomena in ionized gases, whereas Bisek’s results are based on the MHD equations. For hypersonic flow around a hemisphere, the standoff distance increases with increasing Knudsen number when rarefied effects are significant [32].
Figure 8 presents the surface pressure and heat flux distributions for both $Ar$ and $Ar^{+}$. Figure 8(a) illustrates that the surface pressure of $Ar$ increases near the stagnation region and decreases in the shoulder region. A similar trend in surface pressure is observed for $Ar^{+}$. Figure 8(c) shows a decrease in heat flux near the stagnation region for $Ar$, attributable to an enlarged standoff distance leading to a smaller temperature gradient. In contrast, the heat flux of $Ar^{+}$ increases in the stagnation region, potentially due to larger kinetic energy caused by the magnetic field. Given that the mole fraction of $Ar^{+}$ is only 0.00623, the overall surface pressure and heat flux are predominantly governed by $Ar$. The observed increase in stagnation pressure and decrease in heat flux on the hemisphere’s surface are consistent with most established literature in this field [3, 7, 8].
Figure 5: Contour plots of mass density without external magnetic field (upper) and with external magnetic field (lower) in the figures, where (a) ArAr (b) $Ar^{+}$ (c) $e^{-}$ and Knudsen number of ArAr and x-velocity of (d) ArAr (e) $Ar^{+}$ (f) $e^{-}$ and temperature of (g) ArAr (h) $Ar^{+}$ (i) $e^{-}$ at $B_{max}=0.447T$.
Figure 6: Contour plots of (a) electric field magnitude $|\boldsymbol{E}|$ (b) current density magnitude $\boldsymbol{J}$ (c) Lorentz force magnitude $|\boldsymbol{F}|$ at $B_{max}=0.447T$. Upper and lower figures refer to the results without and with external magnetic fields.
@@
Figure 7: Changes in shock standoff distance as a function of magnetic field strength for Mach 4.75 pre-ionized argon flow around a hemispherical geometry. [6] (Experimental uncertainties is $\pm 10\%$).
Figure 8: Surface pressure of (a) ArAr (b) $Ar^{+}$ and surface heat flux of (c) ArAr (d) $Ar^{+}$ at $B_{max}=0.447T$.
3.2.2 Control effects under different Knudsen numbers
[…]
Figure 9: Contour plots of the argon atoms temperature field of (a) Kn=0.2 (b) Kn=0.044 (c) Kn=0.008 at $B_{max}=0.447T$. Upper and lower figures refer to the results without and with external magnetic fields.
Figure 9 displays the argon atom temperature field at various Knudsen numbers. The shock front exhibits a sharpening trend as the Knudsen number decreases. The standoff distance increment reduces with increasing Knudsen numbers. As a result, the total heat flux decrement also exhibits a decreasing trend with larger Knudsen numbers as shown in Table 4. This phenomenon is due to the enhanced decoupling between atoms and charged particles at higher Knudsen numbers. As illustrated in Figure 10, the left panel shows the temperature profiles of argon atoms and ions under an external magnetic field at Kn = 0.2. In this regime, the low collision frequency permits significant temperature decoupling between neutral argon atoms and ions. In contrast, the right panel presents results at Kn=0.008, where the elevated collision frequency drives the temperatures of argon atoms and ions towards equilibrium. In summary, increasing the Knudsen number amplifies the decoupling between charged and neutral particles, thereby weakening the influence of electromagnetic control effects.
Knudsen number Heat flux decrement
0.2 7.09%
0.044 12.37%
0.008 13.71%
Table 4: Heat flux decrement with varying Knudsen number.
Figure 10: Temperature profile of argon atom and argon ions along the stagnation line at (a) Kn=0.2 and (b) Kn=0.008.